OpenStudy (pphalke):

The life span for a certain brand of tires is normally distributed, with a mean of 35,000 miles and a standard deviation of 6,000 miles a. what percent of these tires will last more than 29,000 miles?

jhonyy9 (jhonyy9):

please help me i have understaood it right you need calculi how many percent is 29.000 miles of 35.000 miles ? or something in this way ?

jhonyy9 (jhonyy9):

so i think i ve got it 35000 miles ----- 100% 6.000 miles ---- x % -------------------- x = 6000*100/35000 = 6*100/35 = 17,14% so than now you know that this 6000 miles standard deviation mean 17,14% 35000 - 6000 = 29000 and bc. 35000 mean 100% so than you subtract from 100% the 17,14% you will get the right answer on this question of this exercise 100,00- 17,14 --------- 82,86% so from these result that these 29000 miles mean 82,86% the right answer on this question will be 17,14% do you understand it right now ?